Grade 11Algebra

there are six periods in school and five subjects. If one subject gets at least one period.what are arrangements?

Karan Saini
13 Years agoGrade 11

1 Answer

Aman  Bansal
13 Years ago

Dear Karan

P(5,6)ways

=6*5*4*3*2*1=720*5=3600

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert